the length of a rectangle is 4ft more than twice the with. The area of the rectangle is 48ft squared. Find the length and width of the rectangle.
Answer by checkley77(12844)

Area=LW
L=2W+4
48=(2W+4)W
48=2W^2+4W
2W^2+4W-48=0
2(W^2+2W-24)=0
2(W+6)(W-4)=0
W-4=0
W=4 ANS FOR THE WIDTH.
L=2*4+4
L=8+4
L=12 ANS. FOR THE LENGTH.
PROOF:
48=12*4
48=48